A school with 125 teachers and 4650 students wants to maintain a student-to-teacher ratio of 30:1. How many teachers do they need to hire in order to meet that ratio?
Found 2 solutions by courtney.dee, Alan3354:
Answer by courtney.dee(1)

the Answer would be 155
155*30 = 4650

Answer by Alan3354(69443)

4650/30 = 155
155-125 = 30
They have to hire 30 additional teachers.